# Graph-Informed Neural Networks for Sparse Grid-Based Discontinuity Detectors

### v1

@@ Abstract
In this paper, we present a novel approach for detecting the discontinuity interfaces of a discontinuous function. This approach leverages Graph-Informed Neural Networks (GINNs) and sparse grids to address discontinuity detection also in domains of dimension larger than 3. GINNs, trained to identify troubled points on sparse grids, exploit graph structures built on the grids to achieve efficient and accurate discontinuity detection performances. We also introduce a recursive algorithm for general sparse grid-based detectors, characterized by convergence properties and easy applicability.
Numerical experiments on functions with dimensions $n=2$ and $n=4$ demonstrate the efficiency and robust generalization of GINNs in detecting discontinuity interfaces. Notably, the trained GINNs offer portability and versatility, allowing integration into various algorithms and sharing among users.
Keywords: Discontinuous functions; Sparse grids; Deep learning; Graph neural networks; Discontinuity interface detection.
MSC: 68T07, 03D32, 65D40
@@ 1 Introduction
Detecting discontinuity interfaces of discontinuous functions is a challenging task with significant implications across various scientific and engineering applications. Identifying these interfaces is particularly critical for functions with a high-dimensional domain, as their discontinuities can significantly influence the behavior of numerical methods and simulations; for example, within the realm of uncertainty quantification, where the smoothness of the target function plays a fundamental role in the use of stochastic collocation methods. Specifically, the knowledge of discontinuity interfaces enables the partitioning of the function domain into regions of smoothness, a crucial factor in improving the performance of numerical methods (e.g., see [17]). Other examples of discontinuity detection applications include signal processing, nonlinear partial differential equation (PDE) simulations, investigations of phase transitions in physical systems [14], and change-point analyses in geology or biology, to name a few [30].
@@
The central objective of most discontinuity detection methods is to identify the position of discontinuities in the function domain using function evaluations on sets of points. Over the last few decades, progresses has been made in discontinuity detection, leading to the development of various algorithms. Notable works, such as [3, 2, 16, 35], have introduced significant contributions in this field. In particular, [3] introduced a polynomial annihilation edge detection method designed for piece-wise smooth functions with low-dimensional domains ($n\leq 2$). This method identifies discontinuous interfaces by reconstructing jump functions based on a set of function evaluations. Starting from these results, [2] extended the approach to higher dimensions, applying the detection method for each input dimension within a generalized polynomial chaos approximation of the target function. However, in [2], the curse of dimensionality restricts considerably the practical applicability of the method. A significant advance in addressing the curse of dimensionality was made in [16], where sparse grids (see [24, 7, 21]) were utilized to create an adaptive method. This approach increased the method's potential for use in higher dimensions. On the other hand, few years later, [35] presented a unique approach based on the approximation of hypersurfaces representing the discontinuity interfaces using hyper-spherical coordinates. This method is suitable for large domain dimensions but was primarily designed for detecting a single interface, assuming it satisfies star-convexity assumptions. Concerning the discontinuity detection problem, other techniques include wavelet-based methods [6, 25], filter-based algorithms [15, 31], troubled cell detection [10, 28].
@@
In recent years, Neural Networks (NNs) have been successfully applied also to this problem. For example, in [9], a novel approach emerged, aimed at constructing discontinuous NNs capable of approximating discontinuous functions incorporating trainable discontinuity jump parameters, and enabling the model to learn at the same time both the discontinuity interfaces of the target function and the function itself. This method offered several advantages, with few restrictions on its applicability; actually, the main limit is the curse of dimensionality related to the training set cardinality, which can be very large for functions in high dimensional domains.
@@
We believe that [30] describes a very promising direction; then, in this work, we modify and extend the metodology of [30], moving from regular grids (and CNNs) to sparse grids (and NNs based on graphs), in order to tackle the problem in dimensions higher than 3. In particular, we introduce a novel approach that utilizes Graph-Informed Neural Networks (GINNs) [5] trained for detecting the so-called troubled points in a sparse grid used for the function evaluations; in particular, each GINN is based on the adjacency matrix of a graph built on the structure of the chosen sparse grid. In addition, partially inspired by [16], we define a recursive algorithm for general sparse grid-based detectors, and characterized by finite termination and convergergence properties. Then, running this algorithm using as detector a GINN well-trained in dimension $n$, we obtain a cheap and fast discontinuity detection method suitable for any functions with $n$-dimensional domain. Actually, also Multi-Layer Perceptrons (MLPs) can be trained as sparse grid-based detectors but, according to the results in [5], the experiments show that GINNs take advantage of the graph structure built on the sparse grid, returning better results.
Concerning the experiments, we test our method on functions with domains of dimensions $n=2$ but also on a function with domain of dimension $n=4$, in order to prove its efficiency. The results obtained are promising and show very good generalization abilities of the GINNs in detecting discontinuity interfaces even for functions different from the ones used in the training sets. Moreover, we point the attention of the reader to the fact that the trained GINNs for discontinuity detection have the advantage of being portable and versatile; i.e., they can be shared among different users and integrated into new algorithms. For this reason, we make available the trained GINNs used in the two- and four-dimensional experiments for public use and research (see 5.2).
The work is organized as follows. In Section 2, we recall briefly how to build sparse grids based on equispaced collocation knots and we introduce the concept of sparse grid graph. In Section 3, we give the definition of discontinuity detectors and we introduce the algorithms for finding the discontinuity interfaces of a discontinuous function. Then, in Section 4, we describe the procedure for building and training a NN-based discontinuity detector (both MLP and GINN); further details are illustrated in Appendix A, at the end of the paper. In Section 5, we show the results of the NN-based discontinuity detectors for two-dimensional functions (including images) and four-dimensional functions, proving the efficiency and the promising potentialities of the new method (especially if based on GINNs). We end with some conclusions drawn in Section 6.
@@ 2.1 Sparse Grids
A sparse grid $\mathcal{S}$ in $\mathbb{R}^{n}$ is the union of a collection of $n$-dimensional tensor grids, each one obtained by taking the Cartesian product of a collection of sets of knots. In particular, the recipe for building a sparse grid is the following:
@@
    Choose a criterion to generate sets of collocation knots with cardinality defined by a level-to-knots function $m:\mathbb{N}_{+}\rightarrow\mathbb{N}_{+}$. Specifically, for each discretization level $h\in\mathbb{N}_{+}$, generate a set
    $\mathcal{K}_{i,h}=\{k_{1}^{(i)},\ldots,k_{m(h)}^{(i)}\}\subset[\alpha_{i},\beta_{i}]$
    of collocation knots with respect to the $i$-th dimension of the domain, for each $i=1,\ldots,n$.
@@
    Choose a set of multi-indices $\mathcal{I}\subset\mathbb{N}_{+}^{n}$, each one collecting a different combination of discretization levels $h_{1},\ldots,h_{n}$ for the $n$ dimensions of the domain. Then, for each multi-index $\boldsymbol{h}=(h_{1},\ldots,h_{n})\in\mathcal{I}$, we generate the $n$-dimensional tensor grid
@@
    Generate the sparse grid $\mathcal{S}$ as the union of all the tensor grids generated at the previous step, i.e.:
@@
From the recipe above, we observe that the ingredients characterizing the final aspect of the sparse grid are: the criterion for generating the collocation knots paired with a level-to-knots function (item 1), and the set of multi-indices (items 2 and 3). Several choices are possible, according to the specific task. We refer the reader to [7] for a comprehensive work on sparse grids, and to [21] for a MATLAB library for the generation of sparse grids. For our needs, we focus on the following options:
  • Level-to-knot function: we consider the so-called 2-steps and doubling functions [21], here denoted by $m_{1}$ and $m_{2}$, respectively, and defined as follows:
    $m_{1}(h):=2^{h-1}+1\,,$ (1)
    $m_{2}(h):=\begin{cases}1\,,\quad&\text{if }h=1,\\ m_{1}(h)\,,\quad&\text{otherwise}.\end{cases}$ (2)
  • Collocation knots criteria: we consider equispaced collocation knots
    $\mathcal{K}_{i,h}^{[1]}:=\left\{\left.k_{j}^{(i)}=\alpha_{i}+\frac{\beta_{i}-\alpha_{i}}{m(h)-1}(j-1)\,\right|\,j=1,\ldots,m(h)\right\}\,,$ (3)
    and its variant $\mathcal{K}_{i,h}^{[2]}$ designed for working with the doubling level-to-knot function, i.e.:
    $\mathcal{K}_{i,h}^{[2]}:=\begin{cases}\{\ (\alpha_{i}+\beta_{i})/2\ \}\,,\quad&\text{if }m(h)=1,\\ \mathcal{K}_{1,h}^{[1]}\,,\quad&\text{otherwise}.\end{cases}$
  • Multi-indices selection criteria: the criteria for selecting the set of multi-indices $\mathcal{I}$ are typically based on a so-called rule of the approximation function $r:\mathbb{N}_{+}^{n}\rightarrow\mathbb{N}_{+}$ and on a level of approximation value $h^{*}\in\mathbb{N}^{+}$, such that the set $\mathcal{I}$ is defined as
    :=$\mathcal{I}:=\{\boldsymbol{h}\in\mathbb{N}_{+}^{n}\,|\,r(\boldsymbol{h})\leq h^{*}\}\,.$
    The most used rules [7, 4, 21], typically are
    $r_{\text{prod}}(\boldsymbol{h}):=\prod_{i=1}^{n}h_{i}\,,\quad r_{\text{sum}}(\boldsymbol{h}):=\sum_{i=1}^{n}h_{i}\,,\quad\text{and}\quad r_{\max}(\boldsymbol{h}):=\max_{i=1,\ldots,n}h_{i}\,$ (4)
    that generate the multi-indices sets denoted by $\mathcal{I}_{\text{prod}}(h^{*})$, $\mathcal{I}_{\text{sum}}(h^{*})$, and $\mathcal{I}_{\max}(h^{*})$, respectively, for each fixed $h^{*}\in\mathbb{N}^{+}$; in particular, $\mathcal{I}_{\max}$ is the set of multi-indices characterizing regular grids.
In the following, we show a sparse grid example. Starting from $[\alpha_{i},\beta_{i}]=[-1,1]$ for $i=1,2$, using the level-to-knot function $m_{2}$, the equispaced knots $\mathcal{K}^{[2]}_{i,h}$, and the multi-indices set $\mathcal{I}_{\text{sum}}(6)$, we obtain the 2-dimensional sparse grid illustrated in Figure 1. From now on, for simplicity, we will denote as equispaced sparse grid any sparse grid characterized by an equispaced collocation knots criterion.
Figure 1: Example of sparse grid in $\mathbb{R}^{2}$ obtained with $[\alpha_{i},\beta_{i}]=[-1,1]$ for $i=1,2$, level-to-knot function $m_{2}$, equispaced knots $\mathcal{K}^{[2]}_{i,h}$, and multi-indices set $\mathcal{I}_{\text{sum}}(6)$. Black dotted lines depict the boundary of the sparse grid box (see 2.1).
@@ 2.2 Sparse Grid Graphs
The position of the points of a sparse grid is naturally inclined to define a graph. Nonetheless, to the best of authors' knowledge, in literature there are no formal procedures to build a graph from a sparse grid. Therefore, we introduce here a formal definition for a graph associated to a sparse grid in $\mathbb{R}^{n}$. The graph representations of sparse grids is aimed at exploiting the connections between the grid points in the framework of a discontinuity detection task (see later Section 3).
@@
    $\boldsymbol{x}_{i},\boldsymbol{x}_{j}$ are aligned along one of the $n$ dimensions of the $\mathbb{R}^{n}$ space;
@@
    there is no $\boldsymbol{w}\in\mathcal{S}$, $\boldsymbol{w}\neq\boldsymbol{x}_{i},\boldsymbol{x}_{j}$, such that $\boldsymbol{w}$ belongs to the segment $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$;
@@
In a nutshell, we connect two points of the sparse grid aligned along one of the $n$ dimensions if no other grid points exist between them (item (ii), see Figure 2-left); then, we remove all the edges corresponding to line segments $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$ that intersect other line segments $\overline{(\boldsymbol{x}_{p},\boldsymbol{x}_{p})}$ whose length is smaller than or equal to the one of $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$ (item (iii), see Figure 2-right).
Figure 2: Building a sparse grid graph based on the sparse grid $\mathcal{S}$ in Figure 1. Left. Raw graph obtained connecting the aligned and consecutive points of the sparse grid. Right. Final graph obtained removing from the raw one the edges that intersect an edge of length equal or shorter. The edge opacity depends on the edge weights (see 2.4).
From now on, overloading notation, we will identify the line segment $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$ with the corresponding edge $\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$ in the SGG. Then, the edges of a SGG correspond to line segments in $\mathbb{R}^{n}$ with different length. Therefore, it is natural to exploit the length of line segments to assign weights to edges.
Here, we use the strategy of setting a weight inversely proportional to the distance in $\mathbb{R}^{n}$ between the points defining the edges. A weighted version of SGGs will be useful for improving the training of discontinuity detectors based on Graph-Informed NNs.
@@
We remark that the edge weights introduced in (8) are relative quantities, depending on the relative distance of the points in the sparse grid with respect to $\ell$, and $\ell$ is related to the grid refinement level. Therefore, the WSGG of two similar sparse grids only differ for the set of vertices. This property is summarized in the following proposition.
@@
It is easy to observe that if $\mathcal{S}$ is a sparse grid characterized by equispaced collocation knots and by a hypercubic box of edge $\Lambda$ (i.e., $\beta_{1}-\alpha_{1}=\cdots=\beta_{n}-\alpha_{n}=\Lambda$) then, for each edge $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$, there exists $d_{ij}\in\mathbb{N}$ such that the segment length is $\Lambda/2^{d_{ij}}$. Let $h_{\max}>1$ be the maximum discretization level of $\mathcal{S}$ and $m:=m_{1}(h_{\max})-1=2^{h_{\max}-1}$ (see (1) and (2)); then, $d_{ij}\leq h_{\max}-1$, $\ell=\Lambda/m$, and $\omega_{ij}=2^{(d_{ij}-(h_{\max}-1))}\in(0,1\,]$.

### v4

# Graph-Instructed Neural Networks for Sparse Grid-Based Discontinuity Detectors

@@ Abstract
In this paper, we present a novel approach for detecting the discontinuity interfaces of a discontinuous function. This approach leverages Graph-Instructed Neural Networks (GINNs) and sparse grids to address discontinuity detection also in domains of dimension larger than 3. GINNs, trained to identify troubled points on sparse grids, exploit graph structures built on the grids to achieve efficient and accurate discontinuity detection performances. We also introduce a recursive algorithm for general sparse grid-based detectors, characterized by convergence properties and easy applicability. Numerical experiments on functions with dimensions $n=2$ and $n=4$ demonstrate the efficiency and robust generalization properties of GINNs in detecting discontinuity interfaces. Notably, the trained GINNs offer portability and versatility, allowing integration into various algorithms and sharing among users.
keywords:
Discontinuous functions , Sparse grids , Deep learning , Graph neural networks , Discontinuity interface detection
MSC:
[2020] 68T07 , 03D32 , 65D40
journal: arXiv
[…]
[inst1]organization=Department of Mathematical Sciences, Politecnico di Torino,addressline=Corso Duca degli Abruzzi 24, postcode=10129, state=Turin, country=Italy
[…]
[inst3]organization=Gruppo Nazionale per il Calcolo Scientifico INdAM,addressline=Piazzale Aldo Moro 5, postcode=00185, state=Rome, country=Italy
@@ 1 Introduction
Detecting discontinuity interfaces of discontinuous functions is a challenging task with significant implications across various scientific and engineering applications. Identifying these interfaces is particularly critical for functions with a high-dimensional domain, as their discontinuities can significantly influence the behavior of numerical methods and simulations; for example, within the realm of uncertainty quantification, where the smoothness of the target function plays a fundamental role in the use of stochastic collocation methods. Specifically, the knowledge of discontinuity interfaces enables the partitioning of the function domain into regions of smoothness, a crucial factor in improving the performance of numerical methods (e.g., see [1]). Other examples of discontinuity detection applications include signal processing, investigations of phase transitions in physical systems [2], and change-point analyses in geology or biology, to name a few [3].
@@
The central objective of most discontinuity detection methods is to identify the position of discontinuities in the function domain using function evaluations on sets of points. Over the last few decades, progresses have been made in discontinuity detection, leading to the development of various algorithms. Notable works, such as [4, 5, 6, 7], have introduced significant contributions in this field. In particular, [4] introduced a polynomial annihilation edge detection method designed for piece-wise smooth functions with low-dimensional domains ($n\leq 2$). This method identifies discontinuous interfaces by reconstructing jump functions based on a set of function evaluations. Starting from these results, [5] extended the approach to higher dimensions, applying the detection method for each input dimension within a generalized polynomial chaos approximation of the target function. However, in [5], the curse of dimensionality restricts considerably the practical applicability of the method. A significant advance in addressing the curse of dimensionality was made in [6], where sparse grids (see [8, 9, 10]) were utilized to create an adaptive method. This approach increased the method’s potential for use in higher dimensions. On the other hand, few years later, [7] presented a unique approach based on the approximation of hypersurfaces representing the discontinuity interfaces using hyper-spherical coordinates. This method is suitable for large domain dimensions but was primarily designed for detecting a single interface, assuming it satisfies star-convexity assumptions.
Other techniques addressing the discontinuity detection problem include wavelet-based methods [11, 12], filter-based algorithms [13, 14], and troubled cell detection [15, 16, 17, 3]. In the recent work [18], the authors present a method using the so-called null rules, applied and tested on bivariate functions; in particular, this method analyzes the rules’ asymptotic properties, introducing two indicators for function and gradient discontinuities for the detection of points near discontinuities. Moreover, the method presented in [18] is integrated with adaptive approximation techniques based on hierarchical spline spaces for reconstructing surfaces with discontinuities.
@@
In recent years, Neural Networks (NNs) have also been successfully applied to discontinuity detection. For example, in [19] a novel approach emerged, aimed at constructing discontinuous NNs capable of approximating discontinuous functions incorporating trainable discontinuity jump parameters, and enabling the model to simultaneously learn the target function and its discontinuity interface. This method offered several advantages, with few restrictions on its applicability; actually, the main limit is the curse of dimensionality related to the training set cardinality, which can be very large for functions in high dimensional domains.
@@
We believe that [3] describes a very promising direction; in this work, we modify and extend the metodology of [3], moving from regular grids (and CNNs) to sparse grids (and NNs based on graphs), in order to tackle the problem in dimensions higher than 3. In particular, we introduce a novel approach that utilizes Graph-Instructed Neural Networks (GINNs) [26, 27] trained for detecting the so-called troubled points in a sparse grid used for the function evaluations; in particular, each GINN is based on the adjacency matrix of a graph built on the structure of the chosen sparse grid. In addition, partially inspired by [6], we define a recursive algorithm for general sparse grid-based detectors, and characterized by finite termination and convergergence properties. Then, running this algorithm using as detector a GINN well-trained in dimension $n$, we obtain a cheap and fast discontinuity detection method suitable for any function with $n$-dimensional domain. Actually, also Multi-Layer Perceptrons (MLPs) can be trained as sparse grid-based detectors but, according to the results in [26], the experiments show that GINNs take advantage of the graph structure built on the sparse grid, returning better results.
Concerning the experiments, we test our method on functions with domains of dimension $n=2$ but also on a function with domain of dimension $n=4$, in order to prove its efficiency. The results obtained are promising and show very good generalization abilities of the GINNs in detecting discontinuity interfaces even for functions different from the ones used in the training sets. Moreover, we point the attention of the reader to the fact that the trained GINNs for discontinuity detection have the advantage of being portable and versatile; i.e., they can be shared among different users and integrated into new algorithms. For this reason, we make available the trained GINNs used in the two- and four-dimensional experiments for public use and research (see Remark 5.2).
The work is organized as follows. In Section 2, we recall briefly how to build sparse grids based on equispaced collocation knots and we introduce the concept of sparse grid graph. In Section 3, we give the definition of discontinuity detectors and we introduce the algorithms for finding the discontinuity interfaces of a discontinuous function. Then, in Section 4, we describe the procedure for building and training a NN-based discontinuity detector (both MLP and GINN); further details are illustrated in A, at the end of the paper. In Section 5, we show the results of the NN-based discontinuity detectors for two-dimensional functions (including images, for edge-detection applications) and four-dimensional functions, proving the efficiency and the promising potentialities of the new method, especially if based on GINNs. We end with some conclusions drawn in Section 6.
@@ 2.1 Sparse Grids
Sparse grids were originally introduced by Smolyak in [8] for high dimensional quadrature and since then also used for polynomial interpolation and stochastic collocation (e.g., see [28, 29, 30]). While referring the reader to [9] for a comprehensive work on sparse grids, we sketch here the main ideas.
A sparse grid $\mathcal{S}$ in $\mathbb{R}^{n}$ is the union of a collection of $n$-dimensional tensor grids, each one obtained by taking the Cartesian product of univariate sets of knots hyerarchically built through a number of refinement levels. Starting from a univariate distribution of knots, we consider, on the $i$-th interval $[\alpha_{i},\beta_{i}]$, a sequence of set of knots $\{\mathcal{K}_{i,h}\}_{h\geq 1}$, where the set $\mathcal{K}_{i,h+1}$ is a given refinement of the set $\mathcal{K}_{i,h}$, for each $h\geq 1$. Sparse grids are then obtained by collecting suitable $n$-dimensional tensor products of knot sets $\mathcal{K}_{i,h}$. More precisely, a sparse grid is built with the following steps:
@@
    Upon selecting a univariate knot distribution, choose an initial starting grid $\mathcal{K}_{i,1}$ on each interval $[\alpha_{i},\beta_{i}]$; let $m(1)$ denote the number of nodes in $\mathcal{K}_{i,1}$, for all $i=1,\ldots,n$; then, for $h>1$, $h\in\mathbb{N}$, consider a finer grid $\mathcal{K}_{i,h}$ with $m(h)$ nodes, with $m(h+1)\geq m(h)$; the index $h$ represents the refinement level of $\mathcal{K}_{i,h}$.
@@
    Choose a set $\mathcal{I}\subset\mathbb{N}_{+}^{n}$ of multi-indices; for each $\boldsymbol{h}=(h_{1},\ldots,h_{n})\in\mathcal{I}$, let $\mathcal{S}_{\boldsymbol{h}}$ denote the $n$-dimensional grid defined as
@@
    The sparse grid $\mathcal{S}$ is the union of all the tensor grids generated at the previous step:
@@
Several choices are possible for the univariate knot distribution, for the number of nodes $m(h)$ to be introduced on an interval at a given refinement level $h$, and for the set of multi-indices used to build $\mathcal{S}$. We refer the reader to [9] for a comprehensive work on sparse grids, and to [10] for a MATLAB library for the generation of sparse grids; in [10] the parameter $h$ is called discretization level, and the function $m:\mathbb{N}_{+}\rightarrow\mathbb{N}_{+}$ providing $m(h)$ for each $h$ is called level-to-knots function.
Herein, we focus on equispaced nodes, essentially doubling the number of nodes at each refinement, namely
$m(h):=\begin{cases}1\,,\quad&\text{if }h=1,\\ 2^{h-1}+1\,,\quad&\text{otherwise}.\end{cases}$ (1)
It is worth noting that this choice, upon refinement, yields a nested set of nodes.
As far as the selection of multi-indices set is concerned, let $h^{*}\in\mathbb{N}_{+}$ denote a fixed level of approximation; then, the subset can be described as
$\mathcal{I}(h^{*}):=\{\boldsymbol{h}\in\mathbb{N}_{+}^{n}\,|\,r(\boldsymbol{h})\leq h^{*}\}\,.$
where $r:\mathbb{N}_{+}^{n}\rightarrow\mathbb{N}_{+}$ is suitably chosen. Some examples are [9, 31, 10]:
$r_{\text{prod}}(\boldsymbol{h}):=\prod_{i=1}^{n}h_{i}\,,\quad r_{\text{sum}}(\boldsymbol{h}):=\sum_{i=1}^{n}h_{i}\,,\quad r_{\max}(\boldsymbol{h}):=\max_{i=1,\ldots,n}h_{i}\,.$ (2)
Choice $r_{\text{sum}}$ is known as the Total Degree rule, in the framework of polynomial approximation; choice $r_{\max}$ corresponds to a full tensor grid. The multi-index generated by choices in (2) will be denoted by $\mathcal{I}_{\text{prod}}(h^{*})$, $\mathcal{I}_{\text{sum}}(h^{*})$, and $\mathcal{I}_{\max}(h^{*})$, respectively.
In the following, we show a sparse grid example, starting from $[\alpha_{i},\beta_{i}]=[-1,1]$ for $i=1,2$, from univariate equispaced nodes following (1), and using the multi-indices set $\mathcal{I}_{\text{sum}}(6)$; the obtained 2-dimensional sparse grid is depicted in Figure 1. From now on, for simplicity, we will denote as equispaced sparse grid any sparse grid characterized by an initial equispaced nodes distribution.
Figure 1: Example of sparse grid in $\mathbb{R}^{2}$ obtained with $[\alpha_{i},\beta_{i}]=[-1,1]$ for $i=1,2$, level-to-knot function $m$ (see (1)), equispaced knots $\mathcal{K}^{[2]}_{i,h}$, and multi-indices set $\mathcal{I}_{\text{sum}}(6)$. Black dotted lines depict the boundary of the sparse grid box (see Definition 2.1).
@@ 2.2 Sparse Grid Graphs
The position of the points of a sparse grid is naturally inclined to define a graph. Nonetheless, to the best of authors’ knowledge, in literature there are no formal procedures to build a graph from a sparse grid. Therefore, we introduce here a formal definition for a graph associated to a sparse grid in $\mathbb{R}^{n}$. The graph representations of sparse grids is aimed at exploiting the connections between the grid points in the framework of a discontinuity detection task (see later Section 3).
@@
    $\boldsymbol{x}_{i},\boldsymbol{x}_{j}$ are aligned along one of the axes in the $\mathbb{R}^{n}$ space;
@@
    there is no $\boldsymbol{x}\in\mathcal{S}$, $\boldsymbol{x}\neq\boldsymbol{x}_{i},\boldsymbol{x}_{j}$, such that $\boldsymbol{x}$ belongs to the segment $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$;
@@
In a nutshell, we connect two points of the sparse grid aligned along one of the axes if no other grid points exist between them (item (ii), see Figure 2-left); then, we remove all the edges corresponding to line segments $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$ that intersect other line segments $\overline{(\boldsymbol{x}_{p},\boldsymbol{x}_{p})}$ whose length is smaller than or equal to the one of $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$ (item (iii), see Figure 2-right).
Figure 2: Building a sparse grid graph based on the sparse grid $\mathcal{S}$ in Figure 1. Left. Raw graph obtained connecting the aligned and consecutive points of the sparse grid. Right. Final graph obtained removing from the raw one the edges that intersect an edge of length equal or shorter. The edge opacity depends on the edge weights (see Definition 2.4).
From now on, overloading notation, we will identify the line segment $\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}$ with the corresponding edge $\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$ in the SGG. Then, the edges of a SGG correspond to line segments in $\mathbb{R}^{n}$ with possibly different lengths. Therefore, it is natural to exploit the length of line segments to assign weights to edges.
Here, we use the strategy of setting a weight inversely proportional to the distance in $\mathbb{R}^{n}$ between the points defining the edges. A weighted version of SGGs will be useful for improving the training of discontinuity detectors based on Graph-Instructed NNs.
@@
We remark that the edge weights introduced in (6) are relative quantities, depending on the relative distance of the points in the sparse grid with respect to $\ell$, and $\ell$ is related to the grid refinement level. Therefore, the WSGG of two similar sparse grids only differs for the set of vertices. This property is summarized in the following proposition.
@@
It is easily seen that if $\mathcal{S}$ is a sparse grid characterized by equispaced collocation knots and by a hypercubic box of edge $\Lambda$ (i.e., $\beta_{1}-\alpha_{1}=\cdots=\beta_{n}-\alpha_{n}=\Lambda$) then, for each edge $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$, there exists $d_{ij}\in\mathbb{N}$ such that the segment length is $\Lambda/2^{d_{ij}}$. Let $h_{\max}>1$ be the maximum refinement level of $\mathcal{S}$ and $M:=m(h_{\max})-1=2^{h_{\max}-1}$ (see (1)); then, $d_{ij}\leq h_{\max}-1$, $\ell=\Lambda/M$, and $\omega_{ij}=2^{(d_{ij}-(h_{\max}-1))}\in(0,1\,]$.